# Practical X-ray Gastric Cancer Screening Using Refined Stochastic Data Augmentation and Hard Boundary Box Training

### v3

@@ 6.1 R-sGAIA
The proposed R-sGAIA can probabilistically highlight the gastric folds on a pixel-by-pixel basis based on the medical findings, as shown in Figure 5. In this study, we found that the proposed R-sGAIA achieved 6.0 points and 2.7 points better performance in terms of the F1-score than RandAugment on baseline models 1 and 2, respectively. This confirms the importance of probabilistic augmentation based on medical knowledge. In baseline model 2, R-sGAIA further improveed the F1-score by 1.2 points over the previous sGAIA in a stand-alone comparison. This may be because R-sGAIA used validation data to determine $p(e)$ as a continuous function and was thus able to generate many types of enhancement images that better enhanced the gastric folds than sGAIA, which used a subjectively fixed value. The fact that the function of $p(e)$ with respect to the edge strength $e=E(x,y)$ was approximately between 0.1 and 0.9 is believed to have contributed to the performance improvement by allowing more stochastic variation than sGAIA in the very strong and weak edge strength regions, i.e., by contributing to the diversity of the generated images.
@@ 6.2 HBBT
All recent state-of-the-art box object detection models (e.g., YOLOv7 (Wang et al., 2023)) are trained in such a way that the difference between the boxes specified in the training data and the currently detected boxes simultaneously becomes smaller in both their classes (i.e., contents) and locations. In other words, the sum of the class error and position error for each box is the error of the model, and training is performed to minimize them.
The proposed HBBT is a new and effective strategy that also allows learning from healthy control images, which cannot be used to train conventional object detection models, and retrains over-detected boxes by assigning them a unique “hard-sample” class label to suppress future detections. The HBBT strategy of labeling error-prone regions (i.e., hard samples) with the hard sample class corresponds to increasing the class error of the model for these regions and retraining it to reduce false positives. In other words, the HBBT learning algorithm does not focus on the type of object detection model, because from a machine learning perspective, it is rather a strategy to reduce generic errors in these models. Furthermore, the proposed HBBT actively uses negative samples, in this case images of healthy controls, which cannot be used for training by conventional object detection models.
Although the HBBT is a very simple method that recursively learns to eliminate false positives detected on images, including control images, it alone improves precision by 2.3 points and 3.3 points and F1 score by 4.9 points and 2.9 points on baseline models 1 and 2, respectively. The HBBT is a general-purpose learning method that can be combined with any object detection algorithm, not just gastric cancer detection as in this case. Since the final model detection is determined by the bounding box confidence threshold, HBBT reduces not only the number of false positives but also the number of false negatives (middle case in Figure 6).
@@ 6.3 For practical screening
In this diagnostic support system, the response to the diversity in size and shape of the tumor area to be detected depends on the detectability of the object detection model used, and the range of object sizes that can be detected is determined by the hyperparameters (e.g., anchor size and aspect ratio). In the task of cancer detection from gastric X-ray images, the general settings of the object recognition model are not problematic because the images are taken in a way that the size of the tumor in the image does not appear too small to enable the physician to diagnose it visually. However, when dealing with extremely small tumors, it is necessary to set hyperparameters that take their size into account.
Since R-sGAIA is an online data augmentation method and HBBT is a training method, they do not affect the execution time of the inspection. The processing time of 0.5 s per image under current conditions is considered sufficiently practical, and further improvement of the detection model is expected to increase the speed in the future.
The detection of a gastric cancer region from X-ray images is inherently difficult, not only for automatic diagnostic systems but also for experts. Despite these difficulties, our two technological proposals, R-sGAIA and HBBT, have significantly improved the detection performance of gastric cancer and achieved a performance and speed that can actually be used by physicians as a diagnostic aid. Our diagnostic support system will make a concrete contribution to solving the problem of gastric radiography, which has been considered easy to acquire images but difficult to diagnose. In the future, we aim to examine the results more closely with the help of experts to further improve the system further.
@@ 6.4 Comparison with automated diagnostic support methods for other modalities
Many diagnostic support studies have been conducted on endoscopy, which is the most commonly used method for diagnosing gastric cancer. However, as mentioned above, as mentioned above, most of these studies did not or failed to properly separate the training data from the evaluation data, i.e., the training data is effectively leaked, and reported inappropriately high diagnostic performance. Recent studies with proper separation of these data using more sophisticated machine learning methods Shibata et al. (2020); Ahmad et al. (2023) reported a detection performance for gastric cancer of about 0.7 on an image-based F1 score, which is considered the currently feasible diagnostic accuracy.
Owing to its noninvasiveness and whole-body imaging, X-ray computed tomography (CT) is widely used not only for gastric cancer but also for detecting other organs and diseases. Diagnostic studies of gastric cancer for this modality generally use classical CT radiomics as the analytical factor, including the necessary pre-treatment, and depending on the stage, its diagnostic performance is approximately AUC 0.7 – 0.8 Meng et al. (2020); Huang et al. (2022). The application of deep learning technology is mainly for patient prognosis prediction and segmentation of tumor areas, rather than diagnosis of cancer itself. Hao et al. (2022) showed that radiomic features and features extracted by machine learning techniques (i.e., low-dimensional representations) are effective in predicting the prognosis of patients with gastric cancer. It is expected that these machine learning techniques will be used in the future.
The diagnostic performance of these automated diagnostic support methods appears to be lower than the 95% SE of physicians for gastric cancer; however, this is patient-based, and direct comparisons cannot be made. When automated diagnostic methods are introduced in clinical practice, the actual performance will be much closer to the current SE of physicians because multiple outputs can be integrated to make a decision, and this will be a great aid to physician diagnosis. Although our F1 = 0.578 for the diagnostic performance of gastric X-ray is numerically lower than those of other modalities, it is a box-based evaluation. In addition, since the physician makes a diagnosis by referring to multiple images in a gastric X-ray examination, the ability to indicate the ROI for each image makes our method very useful. Our method is also significant because it is the first time that a gastric X-ray image has been used to aid in the diagnosis of gastric cancer.
@@ 6.5 Limitations of the study
In recent years, it has become known that medical images such as MR images contain features unique to each facility due to differences in imaging equipment, protocols, geometry, etc., and that these differences (so-called domain differences) affect the results when machine learning tasks such as classification are applied as is. The gastric X-ray images used in this study were taken only at medical facilities affiliated with Tokai University. The effect of this domain difference may be observed in the analysis of gastric X-ray images taken at different facilities, but this has not been verified.
@@ 7 Conclusions
In this paper, we have proposed an unprecedented, accurate, and fast gastric cancer screening system from gastric X-ray images based on two technical proposals: (1) R-sGAIA and (2) HBBT. The cancer detection performance of the proposed system (R-sGAIA + HBBT on EfficientDet-D7 network), the system achieves a SE of 90.2%, which is higher than physicians (85.5%) with a processing time of approximately 0.5 seconds per image. At this time, 42.5% of the presented candidate box regions contain cancer. In other words, our system is practical in that it can very well limit the areas that physicians need to check. This performance is 5.9 points higher in F1 score than the same network using RandAugment, a state-of-the-art data augmentation method. The diagnosis of gastric cancer using gastric X-ray is a difficult task that requires long experience. We hope that our proposal will help in the diagnosis, leading to early detection and a reduction in the number of deaths in the near future.

### v4

@@ 6.1 R-sGAIA
The proposed R-sGAIA can probabilistically highlight the gastric folds on a pixel-by-pixel basis based on medical findings, as shown in Figure 5. In this study, we found that R-sGAIA achieved an F1-score improvement of 6.0 points and 2.7 points over RandAugment on baseline models 1 and 2, respectively. This highlights the importance of probabilistic augmentation informed by medical knowledge. In baseline model 2, R-sGAIA further improved the F1-score by 1.2 points over the previous sGAIA in a direct comparison. This improvement may be attributed to R-sGAIA’s use of validation data to determine $p(e)$ as a continuous function, enabling it to generate a wider variety of enhancement images that better highlighted the gastric folds compared to sGAIA, which used a subjectively fixed value. The function of $p(e)$ with respect to edge strength $e=E(x,y)$ ranged approximately between 0.1 and 0.9, which likely contributed to the performance improvement. By allowing more stochastic variation in regions with very strong and weak edge strengths, R-sGAIA enhanced the diversity of the generated images, leading to better detection performance.
@@ 6.2 HBBT
Recent state-of-the-art object detection models (e.g., YOLOv7 (Wang et al., 2023)) are trained to minimize the difference between the boxes specified in the training data and the currently detected boxes, both in terms of their classes (i.e., contents) and locations. In other words, the model’s error is the sum of the class error and position error for each box, and training is performed to minimize these errors.
The proposed HBBT introduces a new and effective strategy that allows learning from healthy control images, which cannot be used to train conventional object detection models. It retrains over-detected boxes by assigning them a unique “hard-sample” class label to suppress future false detections. The HBBT strategy increases the class error for error-prone regions (i.e., hard samples) and retrains the model to reduce false positives. In essence, the HBBT learning algorithm is a strategy to reduce generic errors in object detection models, rather than focusing on the specific type of model. Furthermore, HBBT actively incorporates negative samples, such as images of healthy controls, which are typically excluded from training by conventional object detection models.
Although HBBT is a simple method that recursively learns to eliminate false positives detected in images, including control images, it alone improves precision by 2.3 points and 3.3 points and F1-score by 4.9 points and 2.9 points on baseline models 1 and 2, respectively. HBBT is a general-purpose learning method that can be combined with any object detection algorithm, not just for gastric cancer detection, as demonstrated in this study. Since the final model detection is determined by the bounding box confidence threshold, HBBT not only reduces the number of false positives but also helps minimize false negatives (as seen in the middle case in Figure 6).
@@ 6.3 For practical screening
In this diagnostic support system, the ability to handle the diversity in size and shape of the tumor area to be detected depends on the detectability of the object detection model used. The range of object sizes that can be detected is determined by hyperparameters such as anchor size and aspect ratio. In the task of cancer detection from gastric X-ray images, the general settings of the object recognition model are sufficient, as the images are taken in a way that ensures the tumor size is large enough for visual diagnosis by a physician. However, for extremely small tumors, it is necessary to adjust hyperparameters to account for their size.
Since R-sGAIA is an online data augmentation method and HBBT is a training method, they do not impact the execution time of the diagnostic process. The current processing time of 0.5 seconds per image is considered sufficiently practical, and further improvements to the detection model are expected to increase speed in the future.
Detecting gastric cancer regions from X-ray images is inherently challenging, not only for automatic diagnostic systems but also for experts. Despite these challenges, our two technological innovations, R-sGAIA and HBBT, have significantly improved the detection performance of gastric cancer, achieving a level of performance and speed that can be effectively used by physicians as a diagnostic aid. Our diagnostic support system offers a concrete solution to the challenges of gastric radiography, which has been easy to acquire images but difficult to diagnose. In the future, we aim to work closely with experts to further refine and improve the system.
@@ 6.4 Comparison with automated diagnostic support methods for other modalities
Many diagnostic support studies have focused on endoscopy, the most commonly used method for diagnosing gastric cancer. However, as mentioned earlier, most of these studies did not properly separate the training data from the evaluation data, leading to data leakage and artificially inflated diagnostic performance. Recent studies that have addressed this issue by using more sophisticated machine learning methods (Shibata et al., 2020; Ahmad et al., 2023) have reported detection performance for gastric cancer with an image-based F1 score of approximately 0.7, which is considered the current standard for diagnostic accuracy.
Due to its noninvasiveness and ability to image the entire body, X-ray computed tomography (CT) is widely used for detecting not only gastric cancer but also other organs and diseases. Diagnostic studies of gastric cancer using CT typically rely on classical radiomics, including necessary pre-treatment steps. Depending on the cancer stage, diagnostic performance usually achieves an AUC of 0.7 to 0.8 (Meng et al., 2020; Huang et al., 2022). The application of deep learning in this context primarily focuses on predicting patient prognosis and segmenting tumor areas rather than directly diagnosing cancer. Hao et al. (2022) demonstrated that radiomic features and machine learning-extracted features (i.e., low-dimensional representations) are effective in predicting the prognosis of gastric cancer patients, indicating that these techniques will become more prevalent in the future.
While the diagnostic performance of these automated support methods may appear lower than the 95% SE reported by physicians for gastric cancer, it is essential to note that this comparison is patient-based, making direct comparisons challenging. As automated diagnostic methods are introduced into clinical practice, their actual performance is likely to approach the current SE of physicians, particularly when multiple outputs are integrated to inform a final decision, providing significant support to physicians. Although our F1 score of 0.578 for the diagnostic performance of gastric X-ray images may seem lower than that of other modalities, it is important to recognize that this is a box-based evaluation. Additionally, since physicians diagnose gastric cancer by reviewing multiple images during a gastric X-ray examination, our method’s ability to highlight the ROI in each image makes it highly valuable. Furthermore, our approach is significant as it represents the first use of gastric X-ray images to aid in the diagnosis of gastric cancer.
@@ 6.5 Limitations of the study
In recent years, it has become apparent that medical images, such as MR images, contain features unique to each facility due to differences in imaging equipment, protocols, geometry, and other factors. These so-called domain differences can significantly impact the results of machine learning tasks, such as classification, when applied without adjustment. The gastric X-ray images used in this study were exclusively taken at medical facilities affiliated with Tokai University. While domain differences may affect the analysis of gastric X-ray images obtained from different facilities, this has not been verified in this study.
@@ 7 Conclusions
In this paper, we have proposed an unprecedented, accurate, and fast gastric cancer screening system from gastric X-ray images, built on two key innovations: (1) R-sGAIA and (2) HBBT. The proposed system (R-sGAIA + HBBT on the EfficientDet-D7 network) achieves a SE of 90.2%, surpassing that of physicians (85.5%), with a processing time of approximately 0.5 seconds per image. Notably, 42.5% of the detected candidate box regions contain cancer, meaning our system effectively narrows down the areas that physicians need to examine. This system also achieves an F1 score 5.9 points higher than the same network using RandAugment, a state-of-the-art data augmentation method. Diagnosing gastric cancer using gastric X-rays is a challenging task that requires extensive experience. We hope that our proposal will assist in diagnosis, leading to early detection and a reduction in the number of deaths in the near future.